OpenStudy (anonymous):

In a student survey, 118 students indicated that they speak Spanish, 31 students indicated that they speak French, 13 students indicated that they speak both Spanish and French, and 127 students indicated that they speak neither. How many students participated in the survey? I added all of these and got 289, but that is incorrect--help??

OpenStudy (anonymous):

don't count the students who speak both twice. It's unlear but I think they are included in the French only and Spanish only categories. try 276 and let me know :)

OpenStudy (anonymous):

I tried that too, and 276 is incorrect :/

OpenStudy (anonymous):

N=118+31-13+127=136+127=263

OpenStudy (anonymous):

why would you subtract 13? im guessing that doesn't work either

OpenStudy (anonymous):

becuase the students that speak both Spanish and French are already in the groups that speak Spanish and speak French

OpenStudy (anonymous):

I guess the 13 is already included, but yes very strange problem!!

OpenStudy (anonymous):

sounds about right--just something extra to throw us!! lol thank you Carlos!

OpenStudy (anonymous):

oh. i see why